OpenStudy (anonymous):

Could someone pls help me with this Q. "The line 2y-x+5=0 is a tangent to the curve y=bx+ax^0.5 at point (1,-2). Find the value of a and b.

terenzreignz (terenzreignz):

Calculus, then? Could you find the derivative of \[\Large y = bx + ax^{0.5}\]?

OpenStudy (anonymous):

U mean dy/dx= b+0.5a?

terenzreignz (terenzreignz):

I do mean dy/dx, but there's a little error in your differentiation. :)

OpenStudy (anonymous):

i dont see the error

terenzreignz (terenzreignz):

:) Differentiate this part again: \[\Large y = bx + \color{red}{ax^{0.5}}\] and keep in mind that \[\Large \frac{d}{dx}ax^k = kax^{\color{red}{k-1}}\]

OpenStudy (anonymous):

Thankyou!

terenzreignz (terenzreignz):

You can do it from here, then?

OpenStudy (anonymous):

i think so

terenzreignz (terenzreignz):

Okay then :)